An angle is 30 degrees less than the measure of its supplement. Find the measure of the supplement.
Answer by Theo(13342)

x = angle
180 - x = supplement
x = (180-x) - 30
simplify.
x = 180 - x - 30
add x to both sides
2x = 180 - 30
simplify.
2x = 150
divide both side by 2.
x = 75
angle = 75
supplement = 180 - 75 = 105
75 is 30 less than 105.
solution is that the measure of the supplement is 105.